OpenStudy (anonymous):

What is the acceleration at time t=3 of a particle whose position at time t is given by s(t)=3sqrt(x+1)^(1/3) Note: The square root is the third root and it is being multiplied by another 3 that's the 3 in front.. so, it's 3 * third root(x+1)

OpenStudy (anonymous):

instead of x, it is t?

OpenStudy (anonymous):

\[s(3)=3(\sqrt[3]{x+1})^{1/3}=3\sqrt[9]{4}\]

OpenStudy (phi):

\[s(t)= 3(t+1)^{\frac{1}{3}}\]

OpenStudy (anonymous):

cube root of a cube root? or what?

OpenStudy (anonymous):

it looks like rickjbr but without the 1/3... sorry

OpenStudy (anonymous):

\[3\sqrt[3]{4}\]

OpenStudy (anonymous):

3+1=4 lol

OpenStudy (phi):

replace the cube root with the equivalent to the 1/3 power. take the 2nd derivative of s(t) to find the acceleration

OpenStudy (anonymous):

oh that is the position...i'm retarded

OpenStudy (anonymous):

is the answer -0.333?

OpenStudy (phi):

you mean (t+1) not 4. plug in t=3 after the 2nd derivative

OpenStudy (anonymous):

whoops, i'm not thinking clearly, phi is right.

OpenStudy (phi):

you are in a rush to get the answer. Must be close to quitting time.

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

First derivative is (t+1)^(-2/3)

OpenStudy (anonymous):

2nd would be \[\frac{2(t+1)^{-5/3}}{3}\]

OpenStudy (phi):

with a minus sign

OpenStudy (anonymous):

with a minus sign lol, make it negative

OpenStudy (anonymous):

\[\frac{-2}{3(t+1)^{5/3}}\]

OpenStudy (anonymous):

Now you can plug in 3